(Please maintain at least 2 decimal places in your answers)

QD = 15,000 – 500P

QS = -8,000 + 3,000P

  1. Solve for the equilibrium price and quantity. (Hint: at equilibrium, quantity supplied equals quantity demanded.)

  1. Now assume the supply function changes to:

Qs = -9,000 + 3000P

Is this an increase in supply or a decrease in supply? Calculate the new equilibrium price and quantity
Answer #1

Answer

a)

the equilibrium is at Qd=Qs

equating both equations

15000-500P=-8000+3000P

3500P=23000

P=6.57142857=$6.57

Q=15000-500*6.57142857

Q=11714.2857=11714.29

the price is $6.57, and quantity is 11714.29 units

-----------

b)

The constant term decreases mean the supply shifted to a right which shows an increase in supply.

equaling new supply and old demand

15000-500P=-9000+3000P

3500P=24000

P=24000/3500

P=$6.85714286=6.86

Q=15000-500*6.85714286=11571.4286=11571.43

The new quantity is 11571.43 units, and the price is $6.86